A 100w, 220Vbulb is connected to 160V supply the actual power consumed would be?

Grade:12

1 Answers

kaza manohar
6 years ago
given,
power of bulb=100w,voltage of bulb=220v
P=V2/R
100=2202/R
by solving we get R=222
now,the actual power consumed PA=VA2/R (where VA=ACTUAL VOLTAGE)
PA=1602/222=52.89W